(1) In this problem, we will calculate the electric field of a line charge. The line charge is aligned along the x-axis starting at the origin and having a length L. The line has a linear charge density λ. We want to find the electric field at a point P on the x-axis. The point P is located at the point (d,0) where d > L.

(a) What is the electric field of a small charge dq which is one dimension less than our line?

(b) What is the charge dq in terms of the given parameters?

(c) What is the distance between an arbitrary point on our line charge to the point P in terms of given parameters?

(d) What is the unit vector rˆ in terms of given parameters? (e) What is the electric field of this line charge?
Answer #1

a] E = k*dq/(d-x)^2

b] charge dq = λ*dx

c]distance = d-x

d] rˆ = i^ that is unit vector in x direction

e] E = integral k*λ*dx/(d-x)^2

= 2k*λ*[1/(d-L) - 1/d]